OpenStudy (anonymous):

Solve for Z 1/p+1/k=1/z

OpenStudy (chriss):

\[z(\frac{1}{p}+\frac{1}{k})=1\] \[z=\frac{1}{\frac{1}{p}+\frac{1}{k}}\] z=1*(p+k) z=p+k